# When Large Language Models contradict humans? Large Language Models’ Sycophantic Behaviour

### v3

@@ 3.4 Models
To analyze the sycophantic behaviours of state-of-the-art fine-tuned LLMs, we experiment with three groups of models: two from the OpenAI family OpenAI (2023): GPT-3.5 and GPT-4; two forms of the Meta family Touvron et al. (2023): Llama2-chat-7b, Llama2-chat-13b, and Llama2-chat-70b; and two forms of the Mistral family Jiang et al. (2023, 2024): Mistral-7b and Mixtral-8x7b. We will omit "chat" and the letter "b" for the Meta and Mistral families to simplify the discussion. The resulting names will be Llama2-7, -13, -70 and Mistral-7 and Mixtral-8x7. We used both open-source models - the Meta and Mistral families - to make our work more reproducible and closed-source models - the OpenAI family - because they demonstrate outstanding performance in many NLP tasks. In Appendix C we better describe the characteristics of the models and all the parameters adopted for our experiments.
@@ 4 Results & Discussion
Large Language Models (LLMs) fine-tuned via human feedback appear sensitive to user prompts. In fact, the proposed models, although at different scales, seem to follow user beliefs, although these were provided in opposite directions on topics of politics and philosophy as shown in Figure 5 and discussed in Section 4.1. Moreover, they easily fall into the trick of mimicking user mistakes, as discussed in 4.2.
Finally, although previous experiments demonstrate a variety of weaknesses and a lack of robustness and firmness on the part of LLMs, the sycophantic behaviour towards user interactions appears significantly lower in question-answering tasks. In fact, despite misleading hints, the examined models appear self-confident in their choices as deeply analyzed in Section 4.3.
Figure 5: We investigate the tendency of LLMs to repeat user opinions (sycophancy). Using three benchmark beliefs (Section 3.1), we estimate the percentage of model responses in agreement with the users’ point-of-view.
@@ 4.1 Chameleon LLMs
The human point-of-view manifested via users’ beliefs in prompts tended not to be contrasted by LLMs that reveal a chameleon-like attitude. In fact, by systematically proposing a series of prompts on the same topics with different opinions, LLMs generated responses in line with the views expressed by the users, even though these are totally conflicting among them. This can be seen in Figure 6 and in other instances reported in Appendix D.
@@
The results in Figure 5 show that for Political Questions (POLI-Q) and NLP Research Questions (NLP-Q) related topics, the difference in agreement between the GPTs and Llama2-70 is about 4 points on average. Meanwhile, the difference between the GPTs and Llama2-13 and -7 is about 8 points on average. On the other hand, the Mistral models have an average agreement rate of around 62%. In contrast, in Philosophy-related Question (PHI-Q), we observe the gap only in the Mistrals, which reveal an agreement score of around 72%. Similar but larger scaled values are also present in the Llama models, which is not the case in the GPTs. We believe this is a stronger indication that LLMs produce chameleon-like responses when conversing with humans and expressing their points of view.
@@
Moreover, by analyzing in-family attitudes, it is possible to observe that LLMs with more parameters reveal higher agreement answers than those with fewer parameters; see Llama2-7,-13 and Llama2-70 and both GPTs (in Figure 5) and Appendix D. This phenomenon appears to be directly related to fine-tuning technique, as claimed on a smaller scale with Reinforcement Learning from Human Feedback method in Perez et al. (2022). We observed the same phenomenon in the Mistral models, which, although using a different refinement technique called Direct Performance Optimisation Rafailov et al. (2023), are nevertheless trained to maximize human preferences score.
@@
Downstream of the results discussed and confirmed by experiments on additional models discussed in Appendix D, it is possible to observe that (i) the sycophantic attitudes exhibited by LLMs refined via the RLHF technique are absorbed more by models with high numbers of parameters and (ii) however, refinement techniques using human rewarding policies do not always have these weaknesses in fact models using DPOs have proven to be less sycophantic.
@@
These behaviours revealed by the analyzed models are on topics that do not necessarily have an adequate target response. Therefore, in Section 4.2, we analyzed the behaviour of the LLMs who perform a task related to a potentially misplaced prompt.
4.2 When LLMs fall in mistakes
Even the more robust LLMs seem not to contradict the users’ point of view by generating answers in agreement with it, as revealed in Section 4.1. The results may seem positive, as a model should not be biased under specific topics such as politics, but there may also be weaknesses. Indeed, prioritizing the human point of view by taking the prompt as correct could easily lead LLMs into error. One possible weakness can be observed in Figure 3 on one instance of the Non-Contradiction benchmark and in larger experiments in Figure 6. It appears that the LLMs focused primarily on the performance of the task required from the user rather than pointing out the errors in the prompt. In fact, by systematically asking for a text description written by purposely mistaken authors, we observed the tendency of the LLMs to solve the task without actually addressing the truthfulness of the question (see, for instance, Figure 3).
However, although the general results show a tendency to mimic user errors, the fine-grained results showed that the phenomenon is significantly less when the error is significant, as discussed in Appendix E. In the fine-grained analysis, we discern between especially erroneous prompts with entities related to real poets and several public personages. The results in Figure E demonstrate that the percentage of non-contradiction is higher when the entities are poets; instead, it is lower when the entities are public personages (experimentation setting detailed in Appendix E).
In conclusion, we show that although the models examined tend to satisfy the users’ requests and viewpoints by providing satisfactory answers to the speaker and often without emphasizing possible errors in the prompt, the need remains to examine limited contexts such as those present in question-answering benchmarks. Hence, to study this latter aspect, we continue the analysis in Section 4.3
@@ 4.2 When LLMs fall in mistakes
Figure 6: We investigate the agreement rate with user mistakes in our benchmark (Section 3.2). The considered LLMs tend to mimic human mistakes also when faced with actual error (see Figure 3).
@@ 4.3 Behind Self-confidence lies Robustness?
LLMs seem to be Self-confident in their choices, particularly in question-answering tasks where there is little space for users’ point-of-view and perspectives. However, there are some exceptions, as in Figure 1, where it is possible to observe that GPT-3.5 disagrees with hints in input prompts when they are incorrect, but this is not always true, as Llama2-70 and Mixtral seem to follow the misleading hint.
@@
The best-performing LLMs, i.e., those with higher accuracy values (blue bars in Figure 4), appear not to follow users’ misleading hints and significantly improve performance when the hints are correct. This phenomenon is present mainly in the four multiple-choices question-answering tasks on Llama-2-70, Mixtral, and both GPTs (first four plots in Figure 4), while the gap is not in the same scale in the math-word problem tasks (GSM8K and MultiArith in Figure 4). On the other hand, the models with fewer parameters, i.e., Llama2-7 and Mistral-7, not only have a lower baseline performance than the other models but also seem to follow the users’ hints on all task types to a greater scope (red bars in Figure 4). As can be observed, the rate of Self-confidence does not always have the same performance. Therefore, we heightened the analysis by exemplifying the hints in the different task types and performing a more accurate analysis of the motivations behind following bad hints.
@@
LLMs are sensitive to misleading hints. The results obtained in GSM8K and MultiArith show high agreement rates in misleading hints (red bars Figure 4). Therefore, we repeated the experiments, proposing different kinds of prompts.
Hence, as described and discussed in Appendix G, highly misleading hints do not seem to have the previously discussed effects. Instead, it appears that LLMs produce radically different outputs that contradict the hints provided by users, as shown in Figure G.
Merely as happened in the experiments discussed in Section 3.1, prompts containing misleading information but very close to the target domain (e.g., in the Non-contradiction task, poets close to the actual writer, and in these tasks, numbers very close to the target numbers) raised the bar and the generalization challenges of the LLMs by promoting causal generations that tended to meet and mimic the input prompt.
@@
However, the models with fewer parameters underperform those with more parameters, both in benchmarking with the original prompt and versions with correct and incorrect hints. However, we reproduced the experiments on a limited subset to observe the behaviors in instances that are generally classified correctly, as described in Appendix F.
Figure 7 (discussed in Appendix F) shows that the agreement rates with incorrect hints drop significantly when prompts are altered for which the models generate the correct answer. This result indicates that, although the lower-performing LLMs have been shown to follow prompts with mistakes in the overall experiments, the underlying motivations could be related to poor performances on original tasks on particular subsets of instances.
@@ 5 Related Works
Although human feedback has proven to be an excellent component for refining the interaction between the user and the Large Language Model (LLM), this method can bring some adverse effects, such as sycophancy. In particular, seems that the mechanism of Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2023) stimulates the model to consider the users’ opinion Perez et al. (2022) as a as right truth. However, if the model prefers the users’ opinion over the correct answer, regardless of whether it is correct, it brings robustness and reliability issues.
The initial attitudes related to the input prompt were highlighted by Zhao et al. (2021), who emphasized the propensity of LLMs to provide responses related to the input or commonly present in the pre-training dataset. Building upon this statement, Lu et al. (2022) demonstrated how the specific arrangement of examples can vary the models’ performance from state-of-the-art to random-guessing performance. In a similar way, Turpin et al. (2023) discovered that in a chain-of-thought context Wei et al. (2023a), language models can be easily influenced toward specific responses. Moreover, they showed the presence of high bias factors due to prompt sensitivity.
The sensitivity of the prompt and interactions with users seem to be pivotal points for the study of the resilience of LLMs. Perez et al. (2022), by introducing the concept of sycophancy, showed the behaviours of these models not to contradict human ideas and points of view, in particular, embedded in the prompt. Wei et al. (2023b) proposed a data-level intervention to avoid LLMs’ sycophantic behaviours. Finally, Sharma et al. (2023) adopted the experiments proposed by Wei et al. (2023b) by extending the models under investigation and proposing further data to understand the weight of RLHF in sycophantic behaviours.
In this paper, we propose a comprehensive analysis of the attitudes of LLMs by proposing systematic interventions by influencing prompts with misleading hints and opposite points of view. In particular, our contributions are as follows:
  • We discuss different types of tasks where we probe and analyze the sycophantic behaviour of LLMs via a systematic series of interventions. Hence, starting from existing resources, we extend them by instilling human-influenced beliefs and real or misleading hints.
  • We provide a robust analysis of different LLMs by examining the impact of human feedback-based refinement techniques on their behaviour.
  • We show that LLMs tend to follow the views expressed by the user. However, they do not seem easily corruptible in scenarios where the choice of response is strict.
@@ 6 Conclusion
This paper highlights a critical aspect of Large Language Models and their suggestibility to sycophantic behaviour. While Large Language Models (LLMs) have shown outstanding abilities in solving complex tasks and aligning with human evaluations, this adaptability also introduces a tendency to generate responses that may align more with users’ beliefs rather than factual accuracy. We discern among different scenarios that could induce LLMs to have sycophantic behaviour by proposing different interventions for several tasks. From downstream results, it is possible to observe that LLMs exhibit sycophantic behaviour and agree with user beliefs, especially in situations involving subjective opinions or when factual contradictions are expected. At the same time, these attitudes are significantly less pronounced in objective decision-making scenarios. Although this last conclusion in favor of the robustness of LLMs seems reassuring, the overall analysis highlights the partial robustness of these models and raises critical weaknesses in reliability in critical decision-making scenarios.
Appendix A Appendix
@@
Appendix B Dataset Construction
To evaluate the Self-confidence of LLMs in standard question-answering and math word problem benchmarks, we proposed an intervention at the prompt level, as introduced in Section 2.1. However, we divided our intervention by task type. Regarding question-answering tasks with multiple-choice questions, we constructed the hints in two ways: (i) inserting the target choice and (ii) inserting the first non-target choice present among the possibilities in the hints. In the opposite case, i.e., math word problems, where the target value is numeric, we constructed the hints in two ways: (i) inserting the target value and (ii) inserting the target value perturbed by either adding or subtracting the value one randomly. In this way, we avoided potential biases related to choosing corrupted numbers. In Appendix G, we will study this phenomenon more extensively and the downstream impact.
Appendix C Model and Hyperparameters
In our experimental setting, as introduced in Sec. 3.4, we propose different LLMs:

### v4

@@ 3.4 Models
To analyse the sycophantic behaviours of state-of-the-art fine-tuned LLMs, we experiment with three groups of models: two from the OpenAI family OpenAI (2023): GPT-3.5 and GPT-4; two forms of the Meta family Touvron et al. (2023): Llama2-chat-7b, Llama2-chat-13b, and Llama2-chat-70b; and two forms of the Mistral family Jiang et al. (2023, 2024): Mistral-7b and Mixtral-8x7b. To simplify the discussion, we omit "chat" and the letter "b" for the Meta and Mistral families. The resulting names will be Llama2-7, -13, -70, Mistral-7, and Mixtral-8x7. We used both open-source models - the Meta and Mistral families - to make our work more reproducible and closed-source models - the OpenAI family - because they demonstrate outstanding performance in many NLP tasks. In Appendix D, we better describe the characteristics of the models and all the parameters adopted for our experiments.
@@ 4 Results & Discussion
Large Language Models (LLMs) fine-tuned via human feedback appear sensitive to user prompts. The proposed models, although at different scales, seem to follow user beliefs, even though these were provided in opposite directions on topics of politics and philosophy, as discussed in §4.1. In addition, they easily fall into the trick of mimicking user mistakes, as demonstrated in §4.2.
Finally, although previous experiments argue a variety of weaknesses and a lack of robustness and firmness on the part of LLMs, the sycophantic behaviour towards user interactions appears significantly lower in question-answering tasks. In fact, despite misleading hints, the examined models appear self-confident in their choices, as deeply analysed in §4.3.
Figure 4: We investigate the tendency of LLMs to repeat user opinions (sycophancy). Using three benchmark beliefs (§3.1), we estimate the percentage of model responses in agreement with the users’ point-of-view.
@@ 4.1 Chameleon LLMs
The human point-of-view manifested via users’ beliefs in prompts tends not to be contrasted by LLMs that reveal a chameleon-like attitude. By systematically proposing a series of prompts on the same topics with different opinions, LLMs generate responses that align with the views expressed by the users, even when these are completely contradictory. This can be seen in Figure 5 and scaled on different models as reported in Appendix E.
@@
The results in Figure 4 show that for Political Questions (POLI-Q) and NLP Research Questions (NLP-Q) related topics, the difference in agreement between the GPTs and Llama2-70 is about 4 points, and with Llama2-13 and -7 is about 8 points on average. On the other hand, the Mistral models have an average agreement rate of around 62%. In contrast, in Philosophy-related Question (PHI-Q), we observe the gap only in the Mistrals, which reveal an agreement score of around 72%. Similar but larger scaled values are also present in the Llama models, which is not the case in the GPTs. We believe this is a stronger indication that LLMs produce chameleon-like responses when conversing with humans and expressing their points of view.
@@
Moreover, by analysing in-family attitudes, it is possible to observe that LLMs with more parameters reveal higher agreement answers than those with fewer parameters; see Llama2-7,-13, Llama2-70 and both GPTs (in Figure 4) and Appendix E. This phenomenon appears to be directly related to fine-tuning technique, as claimed on a smaller scale with Reinforcement Learning from Human Feedback method in Perez et al. (2022). We observe the same phenomenon in the Mistral models, which, although using a different refinement technique called Direct Performance Optimisation Rafailov et al. (2023), are nevertheless trained to maximise human preferences score. Finally, to reduce the biased generations, we repeated the experiments, omitting the part about agreement (i.e. ‘Do you agree?’), as shown in Appendix K. The results in Table 10 are comparable with those obtained in the version of the task introduced in §3.1.
@@
Downstream of the results discussed and confirmed by experiments on additional models (Appendix E), it can be observed that the sycophantic attitudes exhibited by LLMs refined via the RLHF technique are absorbed more by models with high numbers of parameters.
@@
However, these behaviours are on topics that do not necessarily have an unquestionable target response. Therefore, in §4.2, we analyse the behaviour of the LLMs that perform a task related to a potentially misplaced prompt.
@@
Figure 5: We investigate the agreement rate with user mistakes in our benchmark (§3.2). The considered LLMs tend to mimic human mistakes also when faced with actual error as in Figure 3.
Figure 6: We examine the Self-confidence and sycophantic behaviours of LLMs in §2.3 in question-answering tasks. We use subsets of four datasets (§3.3). We measure the number of responses in which LLMs agree with the correct (green bar) and incorrect (red bar) hints provided in the prompt.
4.2 When LLMs fall into mistakes
Even the more robust LLMs seem not to contradict the users’ point of view by generating answers in agreement with it, as revealed in §4.1. The results may seem positive, as a model should not be biased under specific topics such as politics, but there may also be weaknesses. Indeed, prioritising the human point of view by taking the prompt as correct could easily lead LLMs into error. One possible weakness can be observed in Figure 3 on one instance of the Non-Contradiction benchmark and in larger experiments in Figure 5. It appears that the LLMs primarily focus on performing the task required by the user rather than pointing out errors in the prompt. By systematically asking for a text description written by purposely mistaken authors, we observe the tendency of the LLMs to solve the task without actually addressing the truthfulness of the question (see, for instance, Figure 3).
However, although the general results tend to mimic user errors, the fine-grained results show that the phenomenon is considerably lower when the error is significant, as discussed in Appendix F. In the fine-grained analysis, we discern between especially erroneous prompts with entities related to real poets and several public personages. The results in Figure F demonstrate that the percentage of non-contradiction is higher when the entities are poets; instead, it is lower when the entities are public persons (experimentation setting detailed in Appendix F). Finally, to complete the analysis in Appendix J, we propose an additional evaluation showing that this phenomenon occurs beyond poetry tasks as well. Specifically, we observed comparable results by proposing a task to summarise biographies belonging to the wrong people. This result shows that the problem is not limited to the poetry task.
In conclusion, we show that although the models examined tend to satisfy the users’ requests and viewpoints by providing satisfactory answers to the speaker and often without emphasizing possible errors in the prompt, the need remains to examine limited contexts such as those present in question-answering benchmarks. To study this latter aspect, we continue the analysis in §4.3
@@ 4.3 Behind Self-confidence lies Robustness?
LLMs seem Self-confident in their choices, particularly in question-answering tasks where there is little space for users’ point-of-view and perspectives. However, there are some exceptions, as in Figure 1, where GPT-3.5 disagrees with hints in prompts when they are incorrect, but this is not always true, as Llama2-70 and Mixtral seem to follow the misleading hint.
@@
The best-performing LLMs, i.e., those with higher accuracy values (blue bars in Figure 6), appear not to follow users’ misleading hints and significantly improve performance when the hints are correct. This phenomenon is present mainly in the four multiple-choices question-answering tasks on Llama-2-70, Mixtral, and both GPTs (first four plots in Figure 6), while the gap is not in the same scale in the mathematical tasks (GSM8K and MultiArith in Figure 6). On the other hand, the models with fewer parameters (Llama2-7 and Mistral-7) not only have a lower baseline performance than the other models but also seem to follow the users’ hints on all task types to a greater scope (red bars in Figure 6). As observed, the self-confidence rate does not always have the same performance. Therefore, we heighten the analysis by exemplifying the hints in the different task types and performing a more accurate analysis of the motivations behind following bad hints.
@@
LLMs are sensitive to misleading hints. The results obtained in GSM8K and MultiArith show high agreement rates in misleading hints (red bars Figure 6). Therefore, we repeat the experiments, proposing different kinds of prompts.
Hence, as described and discussed in Appendix H, highly misleading hints do not seem to have the previously discussed effects. Instead, it appears that LLMs produce radically different outputs that contradict the hints provided by users, as shown in Figure H. Merely as happened in the experiments in §3.1, prompts containing misleading information but very close to the target domain (e.g., in the Non-contradiction task, poets close to the actual writer, and in these tasks, numbers very close to the target numbers) raise the bar and the generalisation challenges of the LLMs by promoting causal generations that tend to meet and mimic the prompt.
@@
However, the models with fewer parameters underperform those with more parameters, both in benchmarking with the original prompt and versions with correct and incorrect hints. However, we reproduce the experiments on a limited subset to observe the behaviours in instances that are generally classified correctly, as described in Appendix G. Figure 8 (discussed in Appendix G) shows that the agreement rates with incorrect hints drop significantly when prompts are altered for which the models generate the correct answer. This result indicates that, although the lower-performing LLMs have been shown to follow prompts with mistakes in the overall experiments, the underlying motivations could be related to poor performances on original tasks on particular subsets of instances.
@@ 5 Related Works
Although human feedback has proven to be an excellent component for refining the interaction between the user and the Large Language Model (LLM), this method can bring some adverse effects, such as sycophancy. In particular, it seems that the mechanism of Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2023) stimulates the model to consider the users’ opinion Perez et al. (2022) as the right truth. However, if the model prefers the users’ opinion over the correct answer, regardless of whether it is correct, it brings robustness and reliability issues. The initial attitudes related to the input prompt were highlighted by Zhao et al. (2021), who emphasised the propensity of LLMs to provide responses related to the input or commonly present in the pre-training dataset. Building upon this statement, Lu et al. (2022) demonstrated how the specific arrangement of examples can vary the models’ performance from state-of-the-art to random-guessing performance. Similarly, Turpin et al. (2023) discovered that in a chain-of-thought context Wei et al. (2023a), language models can be easily influenced toward specific responses. Moreover, they showed the presence of high bias factors due to prompt sensitivity. The sensitivity of the prompt and interactions with users seem to be pivotal points in the study of LLMs’ resilience. Perez et al. (2022), by introducing the concept of sycophancy, showed the behaviours of these models not to contradict human ideas and points of view embedded in the prompt. Wei et al. (2023b) proposed a data-level intervention to avoid LLMs’ sycophantic behaviours. Finally, Sharma et al. (2023) adopted the experiments proposed by Wei et al. (2023b) by extending the models under investigation and proposing further data to understand the weight of RLHF in sycophantic behaviours.
In this paper, we propose a comprehensive analysis of the attitudes of LLMs. Building on previous work (summarised in Table 1), we introduce systematic interventions that influence prompts with misleading hints and opposite points of view. Our contributions are as follows: (i) We discuss different tasks where we probe and analyse the sycophantic behaviour of LLMs via a systematic series of interventions. Hence, starting from existing resources, we extend them by instilling human-influenced beliefs and real or misleading hints. (ii) We provide a robust analysis of different LLMs by examining the impact of human feedback-based refinement techniques on their behaviour. (iii) We show that LLMs tend to follow the views the user expresses, except when the response choice is strict.
@@ 6 Conclusion
This paper highlights a critical aspect of Large Language Models (LLMs) and their suggestibility to sycophantic behaviour. While LLMs have shown outstanding abilities in solving complex tasks and aligning with human evaluations, this adaptability also introduces a tendency to generate responses that may align more with users’ beliefs rather than factual accuracy. We discern scenarios that could induce LLMs to have sycophantic behaviour by proposing different interventions for several tasks. From downstream results, it emerges that LLMs exhibit sycophantic behaviour and agree with user beliefs, especially in contexts involving subjective opinions or when factual contradictions are expected. At the same time, these attitudes are significantly less pronounced in objective decision-making scenarios. However, the overall analysis highlights the partial robustness of these models and raises critical weaknesses in reliability in critical decision-making scenarios.
7 Limitations & Future Works
In this work, we studied the tendencies of LLMs to produce responses in line with users even in the presence of errors or mistakes, a behaviour known as sycophancy. In particular, we analysed this on question-answering benchmarks and observed that the models of the GPT family are very robust and do not get influenced by human-influenced prompts. Although this seemed animating from a stability point of view, it was not confirmed in further analyses. In fact, by systematically asking for opinions in contexts strongly guided by human opinion, the GPTs also did not counter the latter. Finally, we tested the tendency to mimic human errors even in the presence of obvious mistakes. Similarly, the Llama family models and the GPTs showed minor disagreement with prompts specially manipulated to simulate human error.
Appendix A Related Works Schematisation
Work Tasks Models
Perez et al. (2022) POLI-Q, NLP-Q, PHIL-Q encoder-decoder + RLHF (not-released parameters)
Wei et al. (2023b) POLI-Q, NLP-Q, PHIL-Q, Flan, Flan-PaLM (SFT)
MMLU, BBH
Sharma et al. (2023) MMLU, MATH, 4 QA benchmarks Claude 1.3 and 2, GPT-3.5 and -4
non-contradiction task, LLaMA2-70
Ours POLI-Q, NLP-Q, PHIL-Q, 6 QA benchmarks, GPT-3.5, LLaMA2-70, Mixtral8x7
**non-contradiction task (two versions) Mistral-7, LLaMA2-7, LLaMA2-13
***POLI-Q, NLP-Q, PHIL-Q (eval modified) Orca-7,-13, Falcon-7,-40,-180, Yi-6, -34
Table 1: Comparison of tasks and models used in different works. In our *non-contradiction tasks: (i) we have enriched this analysis by including additional characters and showing that the outputs change as the characters change; (ii) we introduced new characters and changed the task to summarising biographies. Moreover, for ***POLI-Q, NLP-Q, and PHIL-Q, we propose original versions and modified versions, changing the prompt and using GPT-4-o as the LLM evaluator.
Appendix B Sycophantic behaviour in math-word problem task
@@
Appendix C Dataset Construction
To evaluate the Self-confidence of LLMs in standard question-answering and math word problem benchmarks, we proposed an intervention at the prompt level, as introduced in Section 2.1. However, we divided our intervention by task type. Regarding question-answering tasks with multiple-choice questions, we constructed the hints in two ways: (i) inserting the target choice and (ii) inserting the first non-target choice present among the possibilities in the hints. In the opposite case, i.e., math word problems, where the target value is numeric, we constructed the hints in two ways: (i) inserting the target value and (ii) inserting the target value perturbed by either adding or subtracting the value one randomly. In this way, we avoided potential biases related to choosing corrupted numbers. In Appendix H, we will study this phenomenon more extensively and the downstream impact.
Appendix D Model and Hyperparameters
In our experimental setting, as introduced in Section 3.4, we propose different LLMs: